In the photoelectric effect, the maximum wavelength for electrons to be emitted
from a surface is 586 nm.

Now light with frequency 7.28 x 1014 Hz is directed on the metal. What is the speed of the emitted
electrons?
Answer #1

we have the work function of the metal W = (6.625*10^-34*3*10^8)/(586*10^-9)

W = 3.391*10^-19 Joules

given the light of a frequency 7.28 x 1014 Hz Energy incident = (6.625*10^-34*7.28 x 10^14) = 4.8*10^-19

By the equation E = W + 0.5*m*v^2

4.8*10^-19 =3.391*10^-19 + 0.5*m*v^2

0.5*9*10^-31*v^2 = 1.432*10^-19

v^2 = (1.432*10^-19)/(0.5*9*10^-31)

v = ((1.432*10^-19)/(0.5*9*10^-31))^0.5

v = 564111.8 m/sec = 564.1118 Km/sec